# RepoGraph: Enhancing AI Software Engineering with Repository-level Code Graph

### v1

@@ 4.2 Experiment results
(i) RepoGraph brings consistent performance gain for all combinations of frameworks and LLM model bases. Specifically, RepoGraph achieves an absolute improvement of $+2.66$ and $+2.34$ in terms of the resolve rate for RAG and Agentless, respectively, which is $99.63\%$ and $8.56\%$ of relative improvement. The notable improvement demonstrates the effectiveness of our RepoGraph in adapting to various scenarios by inducing relevant code context and performing precise code editions. Additionally, our best performance so far by plugging in Agentless, $29.67$, achieves the state-of-the-art performance on the benchmark [^6] https://www.swebench.com/. among all open-source methods.
@@
(iii) Performance gain brought by RepoGraph does not rely on more costs. We also compute and report each method’s average cost and token consumption. By introducing RepoGraph, we manage to reduce the costs associated with managing the entire repository while achieving comparable or even superior performance. As shown in Table 5, the performance improvements achieved by RepoGraph are generally proportional to, or slightly lower than, the corresponding increase in cost. This indicates RepoGraph ’s performance gains are not mainly due to increased token usage.
@@ 5.1 Localization Coverage
+RepoGraph GPT-4o 67.3 $\uparrow$4.6 49.3 $\uparrow$3.0 35.0 $\uparrow$2.7
A crucial step in issue resolution is accurately identifying the correct locations within the code that require modification. Proper localization is essential, as it forms the foundation for generating an effective and accurate patch. Without this step, the quality of the fix may be compromised, leading to incomplete or incorrect solutions. In three granularity, we compute the percentage of problems where the edit locations match the ground truth patch. Namely, file-level, function-level, and line-level. We report that a patch contains the correct location if it edits a superset of all locations in the ground truth patch.
@@ 5.2 Investigation of RepoGraph Variants
In this section, we investigate the efficacy of various combinations of sub-graph retrieval and integration techniques as outlined in Section 3. We explore two sub-graph retrieval variants and two integration methods. Specifically, for sub-graph retrieval, we index $k$-hop ego-graphs where $k$ is set to $1$ and $2$. We limit our exploration to $k$ values up to 2 due to the extensive context required for integration and the potential introduction of noise or irrelevant nodes for $k\geq 3$. For the integration methods, we employ two distinct approaches: (i) directly flattening the textual sub-graph by explicitly detailing the relationships between the search term and its neighboring nodes, and (ii) leveraging an LLM first to summarize the sub-graph in terms of the core modules and salient dependencies, before proceeding with further processing. Detailed implementations of these variants and prompts used can be found in Appendix B.
@@
Table 4 presents the resolve rates for four variants of our method. Notably, while the $2$-hop variant incorporates additional information, directly flattening this information results in the poorest performance, with a resolve rate of $26.00\%$, even lower than the original baseline. In contrast, incorporating summarization via the LLM significantly alleviates context length constraints and enhances information organization, thereby improving performance. For the $1$-hop variant, however, we observed that adding summarization actually degrades performance, reducing the resolve rate to $28.33\%$. We hypothesize that this occurs because the flat $1$-hop ego-graph already contains comprehensive information that fits within the LLM’s context window; thus, summarization may introduce inevitable information loss.
@@ 5.3 Transferibility Test
Table 5: Results on subset of CrossCodeEval with GPT-4o as the backbone LLM.
@@
GPT-4o 10.8 59.4 16.7 48.2
+RepoGraph 28.5 68.3 36.1 61.9
To demonstrate the representational power of RepoGraph for repositories and its transferability to tasks requiring an understanding of repository structures, we conducted experiments using the CrossCodeEval benchmark (Ding et al., 2023). CrossCodeEval is a code completion benchmark designed to emphasize numerous cross-file dependencies. The original dataset consists of 2,665 samples derived from real-world repositories. Due to resource constraints, we randomly selected 500 samples from CrossCodeEval, focusing on problems using Python as the evaluation programming language. For evaluation metrics, we follow the settings in the original paper and measure performance with code match and identifier match metrics, assessing accuracy with exact match (EM), edit similarity (ES), and F1 scores. In our experiments, the search terms are determined to be the function within which the current line is to be completed.
Table 5 demonstrates the results on CrossCodeEval using GPT-4o as the backbone language model. The original GPT-4o model struggles with repository-level tasks, evidenced by an EM score of only $10.8$ for code matching and $16.7$ for identifier matching. These results indicate significant limitations in handling code structure and variable usage in a broader repository context. However, with the integration of the RepoGraph method, there is a substantial improvement across all metrics. Code Match EM improves dramatically to $28.5$, while identifier match EM more than doubles to $36.1$. Similarly, the F1 score for identifier matching jumps from $48.2$ to $61.9$, and the ES for code increases from $59.4$ to $68.3$. These improvements suggest that incorporating repository-level knowledge, as facilitated by RepoGraph, greatly enhances the model’s ability to understand and generate more contextually accurate and semantically consistent code.
@@ 5.4 Error Analysis
Figure 3: Venn diagram of RepoGraph and baseline methods on (a) procedural framework and (b) agent framework. We also plot the error distribution of failing cases against counterparts, e.g., detailed error distribution of $12$ cases RepoGraph succeeds while Agentless fails.
@@
We found that the improvement in agent frameworks is more complementary than procedural frameworks, with larger uniquely resolved cases of $22$ compared with $12$ in procedural frameworks. Together, they make even larger performance of $31.33\%$ and $22.33\%$. The reason could also be attributed to the determinism of procedural frameworks. As a plug-in module, RepoGraph tends to make modifications on existing deterministic processes, resulting in larger overlaps in resolved issues compared with baselines. Agent frameworks, on the other hand, have quite different action distributions with RepoGraph as a plug-in (please refer to Appendix A.2). Therefore, the uniquely resolved cases are more compared with procedural frameworks. For error distributions, contextual misalignment is the most prevalent error type, followed by incorrect localization and regressive fixes for all methods, suggesting that while localization is often correct, the applied solutions may regress or fail to integrate contextually. The phenomenon also echoes with conclusion obtained in Section 5.1. This is intuitive as all the existing methods focus on providing a more comprehensive and desired context for LLMs to solve the task, which fundamentally depends on the power of backbone LLMs. We also found that when integrated with RepoGraph, the proportion of error types “incorrect localization” and “contextual misalignment” largely decreases, indicating that RepoGraph is specifically useful in aggregating the related contexts of the current to-be-fixed issue.
@@ A.2 Agent frameworks
LLM agents have a wide range of applications (Wang et al., 2024a; Yuan et al., 2024). We list all the instructions in every step of agent frameworks. The overall and system instruction is shown in Figure 6.
@@
Appendix B Detailed implementations for each RepoGraph variant
@@ Appendix C Additional Results
C.1 Resolve rate by repository
@@
C.2 Resolve rate by time
We plot the results of the resolve rate in terms of the distribution of releasing time for repositories in Figure 11. Most of the issues were observed in recent years, starting from 2018, with a substantial increase in the total number of issues identified after 2018. In the early years (2012-2016), the number of issues remained relatively low, with both resolved and unresolved counts being minimal. Starting in 2017, there has been a noticeable increase in unresolved issues, with 13 unresolved and only 3 resolved issues. In 2019, the number of resolved and unresolved issues significantly increased, with 19 resolved out of 59 issues. This trend continued to rise until 2020, where 47 issues remained unresolved, and only 17 were resolved, marking the year with the highest number of unresolved issues in the dataset. By 2021 and 2022, the number of unresolved issues slightly decreased, while the resolve rate increased compared to 2020. This suggests an improvement in the system’s ability to address issues in these years. In 2023, although the total number of issues dropped to 30, the proportion of resolved issues remained strong, with 8 resolved out of 22 issues.
Figure 11: Distribution of issues resolved by Agentless +RepoGraph plotted in terms of release time.

### v2

@@ 4.2 Experiment results
(i) RepoGraph brings consistent performance gain for all combinations of frameworks and LLM model bases. Specifically, RepoGraph achieves an absolute improvement of $+2.66$ and $+2.34$ in terms of the resolve rate for RAG and Agentless, respectively, which is $99.63\%$ and $8.56\%$ of relative improvement. The notable improvement demonstrates the effectiveness of our RepoGraph in adapting to various scenarios by inducing relevant code context and performing precise code editings. Additionally, our best performance so far by plugging in Agentless, $29.67$, achieves the state-of-the-art performance on the benchmark [^6] https://www.swebench.com/. We kindly request that reviewers refrain from intentionally checking submitter information on the leaderboard to maintain the integrity of the double-blind review process. among all open-source methods.
@@
(iii) Performance gain brought by RepoGraph does not rely on more costs. We also compute and report each method’s average cost and token consumption. By introducing RepoGraph, we manage to reduce the costs associated with managing the entire repository while achieving comparable or even superior performance. As shown in Table 5, the additional token cost introduced by RepoGraph is justified given the significant performance improvement it enables, particularly when compared to the substantially higher costs incurred by previous methods like SWE-agent. This indicates RepoGraph ’s performance gains are not mainly due to increased token usage.
@@ 5.1 Localization Coverage
    +RepoGraph GPT-4o 67.3 $\uparrow$5.6 49.3 $\uparrow$3.0 35.0 $\uparrow$2.7
A crucial step in issue resolution is accurately identifying the correct locations within the code that require modification. Proper localization is essential, as it forms the foundation for generating an effective and accurate patch. Without this step, the quality of the fix may be compromised, leading to incomplete or incorrect solutions. We compute the percentage of problems where the edit locations match the ground truth patch in three granularity. Namely, file-level, function-level, and line-level. We report that a patch contains the correct location if it edits a superset of all locations in the ground truth patch.
@@ 5.2 Investigation of RepoGraph Variants
In this section, we investigate the efficacy of various combinations of sub-graph retrieval and integration techniques as outlined in Section 3. We explore two sub-graph retrieval variants and two integration methods. Specifically, for sub-graph retrieval, we index $k$-hop ego-graphs where $k$ is set to $1$ and $2$. We limit our exploration to $k$ values up to 2 due to the extensive context required for integration and the potential introduction of noise or irrelevant nodes for $k\geq 3$. For the integration methods, we employ two distinct approaches: (i) directly flattening the textual sub-graph by explicitly detailing the relationships between the search term and its neighboring nodes, and (ii) leveraging an LLM first to summarize the sub-graph in terms of the core modules and salient dependencies, before proceeding with further processing. Detailed implementations of these variants and prompts used can be found in Appendix B.1.
@@
Table 4 presents the resolve rates for four variants of our method. Notably, while the $2$-hop variant incorporates additional information, directly flattening this information results in the poorest performance, with a resolve rate of $26.00\%$, even lower than the original baseline. In contrast, incorporating summarization via the LLM significantly alleviates context length constraints and enhances information organization, thereby improving performance. For the $1$-hop variant, however, we observed that adding summarization actually degrades performance, reducing the resolving rate to $28.33\%$. We hypothesize that this occurs because the flat $1$-hop ego-graph already contains comprehensive information that fits within the LLM’s context window; thus, summarization may introduce inevitable information loss.
@@ 5.3 Transferibility Test
Table 5: Results on the subset of CrossCodeEval with GPT-4o and Deepseek-Coder-V2-Lite-Instruct as the backbone LLMs.
@@
Deepseek-Coder 10.2 57.3 16.6 49.1
+RepoGraph 19.7 67.8 29.3 58.9
GPT-4o 10.5 59.6 16.8 47.9
+RepoGraph 28.7 68.9 36.0 61.3
To demonstrate the representational power of RepoGraph for repositories and its transferability to tasks requiring an understanding of repository structures, we conducted experiments using the CrossCodeEval benchmark (Ding et al., 2023). CrossCodeEval is a multilingual code completion benchmark that is derived from real-world repositories. It is designed to emphasize numerous cross-file dependencies. We focus on problems using Python as the evaluation programming language, leading to 2,665 samples. For evaluation metrics, we follow the settings in the original paper and measure performance with code match and identifier match metrics, assessing accuracy with exact match (EM), edit similarity (ES), and F1 scores. In our experiments, the search terms are determined to be the function within which the current line is to be completed.
Table 5 demonstrates the results on CrossCodeEval using GPT-4o and Deepseek-Coder as the backbone language model. The original LLMs struggle with repository-level tasks, evidenced by an EM score of only $10.8$ for code matching and $16.7$ for identifier matching. These results indicate significant limitations in handling code structure and variable usage in a broader repository context. However, with the integration of the RepoGraph method, there is a substantial improvement across all metrics with both LLMs. Particularly, we find that the improvement brought by RepoGraph is more pronounced when integrated with GPT-4o. A potential reason is that GPT-4o is better at interpreting the additional context information provided for better code reasoning. These improvements suggest that incorporating repository-level knowledge, as facilitated by RepoGraph, greatly enhances the model’s ability to understand and generate more contextually accurate and semantically consistent code.
@@ 5.4 Error Analysis
Figure 3: Venn diagram of RepoGraph and baselines on (a) procedural framework and (b) agent framework on SWE-Bench-Lite. We also plot the error distribution of failing cases against counterparts, e.g., detailed error distribution of $12$ cases RepoGraph succeeds while Agentless fails.
@@
We found that the improvement in agent frameworks is more complementary than procedural frameworks, with larger uniquely resolved cases of $22$ compared with $12$ in procedural frameworks. Together, they make even larger performance of $31.33\%$ and $22.33\%$. The reason could also be attributed to the determinism of procedural frameworks. As a plug-in module, RepoGraph tends to make modifications on existing deterministic processes, resulting in larger overlaps in resolved issues compared with baselines. Agent frameworks, on the other hand, have quite different action distributions with RepoGraph as a plug-in (please refer to Appendix A.2). Therefore, the uniquely resolved cases are more compared with procedural frameworks. For error distributions, contextual misalignment is the most prevalent error type, followed by incorrect localization and regressive fixes for all methods, suggesting that while localization is often correct, the applied solutions may regress or fail to integrate contextually. The phenomenon also echoes the conclusion obtained in Section 5.1. This is intuitive as all the existing methods focus on providing a more comprehensive and desired context for LLMs to solve the task, which fundamentally depends on the power of backbone LLMs. We also found that when integrated with RepoGraph, the proportion of all three error types decreases (specifically for “incorrect localization”), indicating that RepoGraph is specifically useful in aggregating the related contexts of the current to-be-fixed issue.
6 Conclusion and Discussion
This paper introduces RepoGraph, a plug-in module for modern AI software engineering. RepoGraph operates at the code-line level and offers desired navigation through code bases by incorporating and aggregating information at the line level, file level, and repository level through subgraph retrieval of ego-graphs. Extensive experiments on real-world SWE tasks show that RepoGraph significantly boosts the performance of both procedural and agent frameworks. In addition, RepoGraph is proved to be smoothly and effectively transferred to other tasks that require repository-level understanding. Future work could further investigate the generalizability of RepoGraph across different programming languages, frameworks, and developer environments. Real-time execution and feedback could also be explored to further enhance RepoGraph towards a dynamic version for better AI software engineering abilities.
@@ A.2 Agent frameworks
We list all the instructions in every step of agent frameworks. The overall and system instruction is shown in Figure 6.
@@
Appendix B Detailed implementations for RepoGraph
B.1 RepoGraph variant
@@
B.2 RepoGraph components
We first provide a simplified version of Figure 2 in Figure 10, which helps readers quickly catch the overall pipeline of integration with RepoGraph.
Figure 10: A conceptual overview of our proposed RepoGraph.
We also provide an illustration in Figure 11 to explain in detail how we exclude project-independent dependencies mentioned in Section 3.1 Step 2.
Figure 11: A detailed illustration of how we filter the project-dependent relations given a code snippet. Code line colors are illustrated with respect to global and local relations.
@@ Appendix C Additional Results
C.1 Results on other LLMs
We conduct additional experiments on SWE-Bench-Lite with Claude-3.5-Sonnet, and the results are shown in Table 6. We found that with Claude-3.5-Sonnet, the results of baselines are improved a little. This is intuitive as Claude-3.5-Sonnet is designed for better coding tasks. Additionally, integrating RepoGraph with the baseline methods can still boost the performance by a large margin, which further prove the effectiveness and extendability of RepoGraph on different foundational models.
Table 6: Results of RepoGraph on SWE-Bench-Lite test set with Claude-3.5-Sonnet in two research lines, including procedural and agent frameworks. We also report the corresponding cost of each method.
Methods  LLM Accuracy Avg. Cost
resolve # samples patch apply \$ cost # tokens
Procedural frameworks
Agentless GPT-4o 27.33 82 97.33 \$0.34 42,376
Agentless Claude-3.5-Sonnet 27.67 83 94.33 \$0.28 40,984
    +RepoGraph Claude-3.5-Sonnet 30.33 $\uparrow$2.66 91 $\uparrow$8 98.67 $\uparrow$4.34 \$ 0.32 46,238
Agent frameworks
SWE-agent GPT-4o 18.33 55 87.00 \$2.53 498,346
    +RepoGraph GPT-4o 20.33 $\uparrow$2.00 61 $\uparrow$6 90.33 $\uparrow$3.33 \$2.69 518,792
SWE-agent Claude-3.5-Sonnet 23.00 69 86.67 \$1.62 521,208
    +RepoGraph Claude-3.5-Sonnet 25.33 $\uparrow$2.33 76 $\uparrow$7 90.67 $\uparrow$4.00 \$ 1.70 539,942
C.2 Results on localization for both pass and fail cases
In order to highlight the improvement in error localization and correction capability of RepoGraph, we provide a thorough analysis of both pass and fail cases. Results are shown in Table 7. Interestingly, we found that the localization accuracy for failure cases is largely lower than the overall average results. This indicates that RepoGraph fixes the issues by correctly identifying the editing locations, without which the issue is unlikely to be resolved.
Table 7: Percentage of problems for accurate edition localizations for Agentless+RepoGraph with respect to file, function, and line levels. All the numbers are computed from the corresponding generated patches. We also report the accuracy for all pass and failure cases, respectively.
Methods LLM file function line
Agentless +RepoGraph GPT-4o 74.3 54.0 36.7
Pass cases GPT-4o 91.0 79.8 65.2
Failure cases GPT-4o 67.3 43.1 24.6
C.3 Resolve rate by repository
@@
C.4 Resolve rate by time
We plot the results of the resolve rate in terms of the distribution of releasing time for repositories in Figure 13. Most of the issues were observed in recent years, starting from 2018, with a substantial increase in the total number of issues identified after 2018. In the early years (2012-2016), the number of issues remained relatively low, with both resolved and unresolved counts being minimal. Starting from 2017, there is a noticeable increase in unresolved issues, with 13 unresolved and only 3 resolved issues. In 2019, the number of resolved and unresolved issues significantly increased, with 19 resolved out of 59 issues. This trend continued to rise until 2020, where 47 issues remained unresolved, and only 17 were resolved, marking the year with the highest number of unresolved issues in the dataset. By 2021 and 2022, the number of unresolved issues slightly decreased, while the resolve rate increased compared to 2020. This suggests an improvement in the system’s ability to address issues in these years. In 2023, although the total number of issues dropped to 30, the proportion of resolved issues remained strong, with 8 resolved out of 22 issues.
Figure 13: Distribution of issues resolved by Agentless +RepoGraph plotted in terms of release time for repositories.